01/17/2000 OWL shift summary


Subject: 01/17/2000 OWL shift summary
From: Vivian O'Dell (odell@fnal.gov)
Date: Fri Feb 04 2000 - 15:28:13 CST


Shift Summary for owl shift, 17/Jan/2000

On Shift: BobH, Valeri.

Hours of beam available at all intensities: 6.8
Hours of beam available at nominal intensity: 6.8
Hours of E799 data (special runs not included): 4.55
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time:

Stop/start run, tape change, fresh_starts (1.25 hrs), debugging and
tracking
down lost permit (1 hr).

========================================================================
Runs ended this shift:
Run 15548 - in progress - BEAM_799

Tapes brought to FCC this shift:
       none
========================================================================
Accesses during shift:

To track down the lost permit.

========================================================================
Overview of shift:

Good beam all shift with intensity around 9e12. Had problems initializing
new run after fresh start because of lost communication with kumquat
crates, Lost permit in stream 5, crate 31 around 5 am, went on access
and moved L2 for the crate to different discriminator.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
              00:00 Because of ktev1 DDD veto at the end of the
                      previous shift run was stopped and we start
                      shift with drt_fresh_start after changing tapes.
                      Run init hangs on stage3 with no obvious error
                      messages. Logout/login and do fresh start again
                      with the same result. Fresh start again and page
                      Leo and Ed. While we were consulting with Ed, red
                      kumquat button popped out (it took about 15 min
                      after init started). Reboot al kumquat fscc's and
                      do fresh start again.

15548 01:15 Start run.

              03:50 Beam disappears, no messages on TV.

              03:59 Beam is back.

              04:56 Lost permit in stream 5.

              05:15 After multiple kfrends, repeater power cyclings
                      and looking for permit upstairs call MCR and ask
                      for an access.

              05:21 Go on access. As we suspected, permit was lost in
                      crate 31 again for the same reason as yesterday -
                      the crate was not getting L2. Found that 621 NIM
                      has dead channel (which was alive yesterday). Found
                      spare 821 module in neighborhood crate and moved
                      L2 to it. This fixes the problem again.

              05:57 We are out of the access.

              05:59 Beam is back. Ktevana errors - FERA ordinal number
                      mismatch, 3 kfrends (trigger was in error state).

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift



This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:28:17 CST